# EEG Foundation Models: A Critical Review of Current Progress and Future Directions

### v2

@@ 4.3 Model Evaluations
Interpretability analyses: An analysis performed in BrainBERT utilized the intrinsic dimensionality (ID) measure to analyze its task-specific embeddings post-finetuning. ID is a geometric measure used to quantify the minimum number of parameters required to represent the embeddings [68]. Their experiments based on a speech decoding task showed that electrodes within top 10th percentile of IDs mainly fall in the regions of the brain that are associated with speech. Similary, EEGFormer, one of the three models that utilized a vector quantizer to train a neural codebook, analyzed the learned codebook to attribute parts of the input EEG data to downstream task predictions.
Pretraining quality: Some models evaluated the quality of pretraining via qualitative and quantitative assessments. Most qualitative assessments were based on visualization of the data reconstructed by the specific models, which were performed in Mentality, LaBraM, and BrainBERT. Quantitative assessments focused on reconstruction error (e.g., Mentality) and performance on forecasting and imputation tasks (e.g., Brant, FoME).
Ablations: Few models evaluated the contributions of various model components or datasets to task performance via ablations. An experiment performed in Neuro-GPT, which included a convolutional-transformer and a GPT decoder, showed that removing the GPT decoder improved downstream performance significantly. Experiments in Brant evaluated the contributions of the temporal encoding, spatial encoding, and the frequency encoding and demonstrated that while all of those components were valuable to learn EEG representations, the contributions of the temporal encoder were the highest. Similarly, LaBraM showed that the spatial embeddings are crucial for the pretraining to converge as well as to perform well in downstream evaluations after pretraining. Furthermore, LaBraM studied the effect of the codebook during pretraining by removing the neural tokenizer with traditional raw EEG (& frequency) masked reconstruction and shows that it loses performance with complex tasks such as TUEV on downstream. NeuroLM showed that different neural tokenizer training tasks (frequency vs. temporal vs. frequency & temporal reconstruction), can be beneficial for different downstream tasks. Additionally, NeuroLM studied the impact of the duration of model pretraining by fine-tuning different checkpoints, and the results were inconclusive. An experiment in BrainWave performed a data ablation by training separate models with scalp and intracranial EEG and compared the performance against joint scalp-iEEG training. Their results indicated that joint training generally improved downstream performance in all tasks, except one.
@@ 5 Key Takeaways and Research Gaps
This section presents several key observations made across the preceding comparative analyses (Section 4) and highlights the research gaps that remain. We believe that addressing these gaps would make future generations of EEG-FMs more principled in their methodology and trustworthy for scientific and clinical investigations.
@@
(b) Model scaling and performance gains. Model sizes are specific to each study. ‘Sm.’ - smallest variant, ‘Md.’ - intermediate variant, ‘Lg.’ - largest variant.
Figure 4: The impact of self-supervised pretraining and model scaling on task performance. Note that downstream tasks and their performance metrics are different within and across EEG-FM studies.
@@ 5.1 Key Takeaways
Diversity of pretraining data: Several EEG-FMs (LaBraM, NeuroLM, FoME, and BrainWave) leveraged a diverse set of EEG domains spanning clinical, sleep, and task-based BCI. LaBraM’s lead in the TUAB and TUEV evaluations (Figure 3) may have emerged from its higher diversity in pretraining data compared to EEGFormer and BIOT. LaBraM also performed competitively in seizure classification tasks on the MAYO and FNUSA datasets as reported in FoME’s evaluations. Note that LaBraM was pretrained on significantly fewer channel-hours of scalp EEG data and while TUAB and TUEV evaluations are both in-distribution assessments, MAYO-FNUSA evaluations are both intracranial EEG datasets. The mixed use of scalp and intracranial EEG, as done in FoME and BrainWave, can be considered as another form of data diversity. Using data ablations (scalp EEG vs. iEEG), BrainWave demonstrated that joint pretraining (scalp EEG + iEEG) leads to a boost in downstream task performance and transfer performance to unseen data types (electrocardiograms). Overall, the notion of data diversity, perhaps in conjunction with data volume, may influence the performance, generalizability, and transferability of EEG-FMs.
Minimal data preprocessing: Most EEG-FMs perform minimal and simple data preprocessing steps, namely filtering and resampling, to standardize EEGs from various sources. Notably, the removal of noise-related outliers, suppression of EEG artifacts, artifactual EEG components, and site-related harmonization were not explicitly pursued. Moreover, data normalization strategies that produce training-ready samples were not sufficiently described in most studies. It remains unclear whether or how various offline data handling strategies impact EEG-FM pretraining and downstream task performance, particularly with OOD test data.
@@
Limited model evaluations: Task performance after fine-tuning was the primary paradigm of EEG-FM evaluation. However, in half the reviewed studies (BrainBERT, Brant, Mentality, NeuroLM, and FoME), the downstream evaluation datasets were already utilized for FM pretraining, i.e., the evaluations were in-sample. Although several studies conducted external evaluations on unseen datasets, BrainWave is the only study that performed truly OOD task evaluations, i.e., without fine-tuning the model on the evaluation set. Direct model rankings beyond the TUAB and TUEV tasks are difficult to determine due to heterogeneous selections of downstream tasks across most EEG-FMs (see Table 2). Even when considering the TUAB and TUEV tasks, only four out of the ten EEG-FMs can be ranked, as depicted on the y-axis in Figures 3(a) and 3(b), respectively. In addition, linear probing or few-shot evaluations were rarely reported (see Fig. 4(a)). Overall, the number of meaningful EEG classification tasks evaluated in each study ranged from 1 (Mentality) to 12 (BrainWave), with most studies evaluating proposed EEG-FMs on at most 5 tasks. Overall, the universality and robustness of EEG-FMs has not been convincingly demonstrated in most studies, with BrainWave being a notable exception.
@@
Performance of general-purpose time series models: Interestingly, we observed that general time series foundation models, such as TimesNet [150], performed reasonably well on several EEG tasks post-fine-tuning, and sometimes outperformed EEG-FMs (e.g., sleep-stage classification in FoME). Additionally, experiments in BrainWave show that, a time series FM – MOMENT [149] – outperformed EEG-FMs in specific tasks, such as seizure detection. Experiments in Brant show that general time series architectures, such as PatchTST [151] and CoST [131], perform relatively better in some tasks, such as short/long term signal forecasting and imputation, respectively, than some EEG-specific architectures. These findings highlight that certain EEG tasks could be tackled using general time series architectures or FMs without any EEG-specific inductive biases.
Data scaling and task performance: The trend along x-axis of Figures 3(a) and 3(b) suggests that scaling up pretraining data may lead to worse downstream task performance, even with significant model scale up (e.g., LaBraM vs. EEGFormer/NeuroLM). Notably, LaBraM demonstrated that effects of pretraining data scaling on TUAB and TUEV classifications are sharpest under $\sim$1000 hours of data and begin to plateau thereafter. Overall, the evidence for data scaling is weak, if any, based on the limited shared tasks and models evaluated thus far.
Advance over other feature paradigms: In Figure 4(a), we compare progressive feature extraction paradigms (expert features with classical ML, data-driven supervised DL features, self-supervised DL features, and proposed EEG-FMs) on various tasks. In a majority of tasks, fine-tuned EEG-FMs (‘D’) provided at least some improvement, if not drastic, over previous DL baselines and FMs (‘B’ and ‘C’, respectively). Linear probing results reported in EEG-FMs, however, were relatively worse in comparison. Fine-tuned EEG-FMs showed substantial improvements over classical ML models with expert features (‘A’), although such assessments were presented only in four studies.
@@ 5.2 Current Research Gaps
Preprocessing and normalization effects: EEG datasets can require significant offline preprocessing to manage data quality and suppress artifacts. Since current studies perform minimal processing, it is unclear whether or to what extent data outliers and noise impact EEG-FM pretraining and task performance. Even clean EEG datasets require careful consideration of normalization strategies as EEG can vary across channels, subjects, and acquisition sites [6]. The choice of input representation can further complicate decisions related to preprocessing and normalization. As such, there is a need to better understand the impact of these choices on downstream EEG-FM modeling.
@@
Long temporal context and spatial modeling: Slow variations over long timescales can exist in multi-day intracranial EEGs or multi-hour sleep-related EEG recordings [152]. However, current EEG-FMs can only process patterns within sequences of 90 seconds or less. There is a need for solutions that expand the effective context length of EEG-FMs. Moreover, the explicit modeling of spatial or inter-channel relationships and their contributions relative to temporal modeling requires further investigation.
@@
Quality of pretraining strategy: EEG-FM linear probing evaluations reported by some studies have performed significantly worse than fine-tuned versions and other baselines in several instances (see Figure 4(a)). However, some gains can be seen when FMs are fine-tuned on task data compared to fully-supervised training on the same data. This contrasting observation casts doubt on the inherent/intrinsic quality of the representations learned via self-supervision in EEG-FMs. Further investigations into the effects of SSL pretraining on downstream evaluations are needed to fully understand the extent of transferability achieved through SSL. Beyond the SSL strategy itself, the effects of data diversity, data volume, and model scale on the quality of EEG-FM pretraining remain unknown.
@@
Practically relevant evaluations and metrics: Current fine-tuning evaluations are limited in their capacity to assess the real-world practical utility of EEG-FMs. There is a need to adopt evaluation schemes and task metrics that capture the reality of EEG research and clinical use. Zero-shot evaluations are required for off-the-shelf EEG-FM usage for novel, unseen tasks. Few-shot or low-label performance can assess how efficiently EEG-FMs can leverage EEG labels that are typically expensive and laborious to collect. Out-of-distribution performance, especially without fine-tuning, on a known task can help understand EEG-FM robustness to the idiosyncratic cross-subject and cross-site variability of EEG. Finally, application-specific metrics, such as false positives per hour for seizure detection, and comparisons with expert EEG feature baselines can further clarify the real-world utility of EEG-FMs.
@@
Standardized benchmarking tasks: The comparative analysis revealed significant heterogeneity in the tasks selected for EEG-FM evaluation (see Table 2). The lack of common tasks across EEG-FM evaluations makes it challenging to understand the state-of-the-art, and highlights a need to identify a common core set of evaluations for future EEG-FM evaluations. This set must cover multiple task types and include both classifications and regressions, with dense (one label for the entire EEG recording) and sparse (one label for each EEG segment) labels. Additionally, the tasks must be challenging for previous generations of EEG-DL models with ample room for improvement, unlike TUAB, where performance may have already saturated (85-87% accuracy) with traditional approaches [153].
@@
Trustworthy modeling: None of the reviewed EEG-FMs focused on model explainability or interpretability, which remain key requirements for data-driven modeling in high-risk and expert-centric domains such as medicine. Studies that demystify the EEG-FM black box are needed to gain insight into knowledge learned by EEG-FMs (EEG patterns, dependencies, relationships) through pretraining and the practical robustness of their decision-making process for downstream applications. Connections to known patterns of brain physiology or pathology may be necessary to make EEG-FMs trustworthy in the eyes of expert users.
@@ 6 Proposed Future Directions
From a time series perspective, EEG signals are distinct from general signals or sequences due to their unbounded real-valued amplitude, non-linearity[154], non-stationarity[154], and 1/f spectral characteristics[155]. These peculiarities would suggest that foundation modeling for EEG signals requires domain-specific approaches. Indeed, we have cultivated an EEG-centric, domain-sensitive ethos throughout this review. However, counter-intuitively, EEG-FMs have enjoyed some success in drawing from vision and language domains in their adoption of sequence patching and transformer architecture design (transformer [45], vision transformer [62]), masked reconstruction-based SSL (masked autoencoders [124]), and discrete neural codebooks (VQ-VAE [123], BEiT v2 [156]). Moreover, time series FMs pretrained on generic, non-neural time series have performed similarly to EEG-FMs on certain EEG tasks. Therefore, it is likely that future EEG-FMs could benefit from embracing both domain-specific insights and innovations from external data domains. These domains may be semantically distant to EEG, such as vision and natural language, or loosely related, such as non-neural time series, audio, speech, or other biosignals. Below, we outline the future research directions that we believe could support meaningful and sustained progress in the EEG-FM research domain. As illustrated in Figure 5, these recommendations are organized under three broad themes of development spanning benchmarks and software tools, technical advances, and real-world applications. These suggestions build upon the research gaps identified earlier and provide additional guidance for advancing and accelerating EEG-FM research and adoption.
Figure 5: Illustration of future directions along three key developmental axes; benchmarks and tools, technical modeling, and applications. I. Future EEG-FMs can be compared against prevailing feature paradigms and participate in EEG challenges to establish their real-world utility. Frictionless and user-friendly software tools are needed to quickly adopt and experiment with off-the-shelf models. II. Holistic evaluation frameworks that test embedding space semantics, robustness, and transfer efficiency can meaningfully track the state-of-the-art. Advanced representation learning techniques, such as federated or multi-modal learning, can enhance large-scale pretraining. III. Collaborations with domain experts can inspire novel EEG-FM applications. Although current applications are largely discriminative, future applications could be generative in nature.

### v3

@@ 4.3 Model Evaluations
Interpretability analyses: An analysis performed in BrainBERT used the intrinsic dimensionality (ID) measure to assess task-specific embeddings post-finetuning. ID is a geometric measure that quantifies the minimum number of parameters required to represent embeddings [ansuini2019intrinsic]. They showed that, in their model, the electrode ID distribution lies in a lower-dimensional space compared to a randomly initialized model. Furthermore, they showed that electrodes in the top 10th percentile of IDs mainly fall in brain regions associated with the task. Similarly, EEGFormer, one of the three models that utilized a vector quantizer to train a neural codebook, analyzed the learned codebook to attribute parts of the input EEG data to downstream task predictions. In their analysis, codes most likely to predict seizures were mapped back to the EEG traces to visualize where the model focused. These highlighted regions corresponded to known epileptiform patterns, indicating that the model’s decisions aligned with meaningful clinical markers.
Pretraining quality: Some models evaluated pretraining quality through qualitative and quantitative assessments. Most qualitative assessments were based on visualizations of data reconstructed by the specific models, which were run in Mentality, LaBraM, and BrainBERT. Quantitative assessments focused on reconstruction error (e.g., Mentality) and performance on forecasting and imputation tasks (e.g., Brant, FoME).
Ablations: Few models evaluated the contributions of various model components or datasets to task performance via ablations. An experiment in Neuro-GPT, which combined a convolutional transformer and a GPT decoder, showed that removing the GPT decoder significantly improved downstream performance. Experiments in Brant evaluated the contributions of temporal, spatial, and frequency encoders and demonstrated that, while all components were valuable for learning EEG representations, the temporal encoder made the greatest contribution. Additionally, LaBraM showed that spatial embeddings are crucial for pretraining to converge and for downstream performance. Furthermore, LaBraM studied the effect of the codebook during pretraining by replacing the neural tokenizer with traditional time- and frequency-domain masked reconstruction and showed that this degraded performance for complex downstream tasks such as TUEV. NeuroLM showed that different neural tokenizer training tasks (frequency vs. temporal vs. frequency and temporal reconstruction) can be beneficial for different downstream tasks. Additionally, NeuroLM examined the impact of model pretraining duration by fine-tuning different checkpoints, but the results were inconclusive. An experiment in BrainWave performed data ablation by training separate models with scalp and intracranial EEG and compared their performance with joint scalp-iEEG training. Their results indicated that joint training generally improved downstream performance in all tasks, except one.
@@ 5 Key Takeaways and Research Gaps
This section presents several key observations made across the preceding comparative analyses (Section 4) and highlights the research gaps that remain. We believe that addressing these gaps would make future EEG-FMs more principled in their methodology and trustworthy for scientific and clinical investigations.
(a) Impact of progressive learning paradigms (A → D) on downstream EEG classifications.
(b) Model scaling and performance gains.
@@ 5.1 Key Takeaways
Diversity of pretraining data: Several EEG-FMs (LaBraM, NeuroLM, FoME, and BrainWave) leveraged a diverse set of EEG domains spanning clinical, sleep, and task-based BCI. Notwithstanding potential confounders such as preprocessing, data splits, and model adaptation strategies, LaBraM’s lead in the TUAB and TUEV evaluations (Figure 3) may have emerged from its higher diversity in pretraining data, i.e., the number of distinct data sources, compared to EEGFormer and BIOT. LaBraM also performed competitively in seizure detection and events classification tasks on the MAYO and FNUSA datasets as reported in FoME’s evaluations. Note that LaBraM was pretrained on significantly fewer channel-hours of scalp EEG data and while TUAB and TUEV evaluations are both in-distribution assessments, MAYO-FNUSA evaluations are both iEEG datasets. The mixed use of scalp and iEEG, as done in FoME and BrainWave, can be considered as another form of data diversity. Using data ablations (scalp EEG vs. iEEG), BrainWave demonstrated that joint pretraining (scalp EEG + iEEG) boosts downstream task performance and transfer to unseen data types (electrocardiograms). Overall, the notion of data diversity, in conjunction with data volume, may influence the performance, generalizability, and transferability of EEG-FMs.
Minimal data preprocessing: Most EEG-FMs perform minimal and simple data preprocessing steps, namely filtering and resampling, to standardize EEGs from various sources. Notably, the removal of noise-related outliers, suppression of EEG artifacts, and site-related harmonization were not explicitly pursued. Moreover, data normalization strategies that produce training-ready samples were not sufficiently described in most studies. It remains unclear whether or how various offline data handling strategies impact EEG-FM pretraining and downstream task performance, particularly with out-of-distribution test data.
@@
Limited model evaluations: Task performance after fine-tuning was the primary paradigm of EEG-FM evaluation. However, in four of the ten reviewed studies (BrainBERT, Mentality, NeuroLM, and FoME), the downstream evaluation datasets were already used for pretraining, i.e., the evaluations were in-distribution. However, notably, BrainWave and Brant are the only studies to have performed out-of-distribution evaluations using unseen datasets (see Supplementary Table 2). Direct model rankings beyond the TUAB and TUEV tasks are difficult to determine due to heterogeneous selections of downstream tasks across most EEG-FMs (see Table 2). Even when considering the TUAB and TUEV tasks, only four out of the ten EEG-FMs can be ranked, as depicted on the y-axis in Figures 3(a) and 3(b), respectively. In addition, linear probing and few-shot evaluations were rarely reported (see Fig. 4(a) and Supplementary Table 2, respectively). Overall, the number of meaningful EEG classification tasks evaluated in each study ranged from 1 (Mentality) to 12 (BrainWave), with most studies evaluating proposed EEG-FMs on at most 5 tasks. Overall, the universality and robustness of EEG-FMs have not been convincingly demonstrated in most studies, with BrainWave being a notable exception.
@@
Performance of general-purpose time series models: Interestingly, we observed that general time series foundation models, such as TimesNet [wu2022timesnet], performed reasonably well on several EEG tasks after fine-tuning and sometimes outperformed EEG-FMs (e.g., sleep-stage classification in FoME). Additionally, experiments in BrainWave show that a time series FM – MOMENT [goswami2024moment] – outperformed EEG-FMs in specific tasks, such as seizure detection and pathology detection. Experiments in Brant show that general time series architectures, such as PatchTST [nie2022time] and CoST [woo2022cost], perform relatively better on some tasks, such as short- and long-term signal forecasting and imputation, respectively, than some EEG-specific architectures. However, other results reported in BrainWave and FoME suggest that EEG-specific inductive biases and modeling choices (e.g., spatial modeling, multi-scale modeling of intracranial and scalp data, time-frequency representations, diverse clinical characteristics) can indeed help EEG-FMs outperform general time-series FMs (TS-FMs) in EEG-specific tasks. Overall, these findings indicate that TS-FMs may provide moderate value in certain EEG tasks due to efficient model adaptation via fine-tuning or larger-scale general pretraining on time series containing richer and more diverse semantics than EEG.
Data scaling and task performance: The trend along the x-axis of Figures 3(a) and 3(b) suggests that scaling up pretraining data may not necessarily increase downstream task performance, even with significant model scale-up (e.g., LaBraM vs. EEGFormer/NeuroLM). Notably, LaBraM demonstrated that the effects of pretraining data scaling on TUAB and TUEV classifications are sharpest under $\sim$1000 hours of data and begin to plateau thereafter. Overall, the evidence for data scaling is weak, if any, based on the limited shared tasks and models evaluated thus far.
Advance over other feature paradigms: In Figure 4(a), we compare progressive feature paradigms (i.e., expert features, supervised EEG-DL features, SSL baselines, and proposed EEG-FMs) on various tasks. In a majority of tasks, proposed EEG-FMs (‘D’) provided at least some improvement after fine-tuning, if not drastic, over previous DL (‘B’) and self-supervised baselines, including previous EEG-FMs (‘C’). Linear probing results, however, were relatively worse. Fine-tuned EEG-FMs showed substantial improvements over classical machine learning (ML) models with expert features (‘A’), although such assessments were reported in only four studies.
@@ 5.2 Current Research Gaps
Preprocessing and normalization effects: EEG datasets can require significant offline preprocessing to manage data quality and suppress artifacts. Since current studies perform minimal processing, it is unclear whether or to what extent data outliers and noise impact EEG-FM pretraining and task performance. Even clean EEG datasets require careful consideration of normalization strategies as EEG can vary across channels, subjects, and acquisition sites [wagh2022evaluating]. The choice of input representation can further complicate decisions related to preprocessing and normalization. As such, there is a need to better understand the impact of these choices on downstream EEG-FM modeling.
@@
Long temporal context and spatial modeling: Slow variations over long timescales can be observed in multi-day intracranial and scalp EEGs [mivalt2023impedance]. Additionally, there is substantial spatial and temporal variation in EEG across behavioral states (e.g., deep sleep versus wake, eyes open versus eyes closed). However, current EEG-FMs can only process patterns within EEG sequences spanning 90 seconds or less. There is a need for solutions that expand the effective context length of EEG-FMs. Moreover, the explicit modeling of spatial or inter-channel relationships and their contributions relative to temporal modeling remains to be investigated.
@@
Quality of pretraining strategy: EEG-FM linear probing evaluations reported by some studies have performed significantly worse than fine-tuned versions and other baselines in several instances (see Figure 4(a)). However, some gains can be observed when FMs are fine-tuned on task data compared to fully supervised training on the same data. This contrasting observation casts doubt on the quality of the representations learned via self-supervision in EEG-FMs. Further investigations into the effects of SSL pretraining on downstream evaluations are needed to fully understand the extent of transferability achieved through SSL. Beyond the SSL strategy itself, the effects of data diversity, data volume, and model scale on the quality of EEG-FM pretraining remain unknown.
@@
Practically relevant evaluations and metrics: Current fine-tuning evaluations are limited in their ability to assess the practical utility of EEG-FMs in real-world settings. There is a need to adopt evaluation schemes and task metrics that capture the reality of EEG research and clinical use. Evaluations on out-of-distribution data and novel tasks are required to assess the off-the-shelf value of EEG-FMs. Few-shot or low-label performance, when measured in absolute terms rather than percentages, can capture how efficiently EEG-FMs leverage labels that are typically expensive and laborious to collect. Out-of-distribution performance on a known task without fine-tuning can help understand EEG-FM robustness to the idiosyncratic cross-subject and cross-site variability of EEG. Furthermore, application-specific, practically relevant metrics, such as false positives per hour for seizure detection, and comparisons with expert features can clarify the real-world utility of EEG-FMs. Finally, evaluations must account for the unbalanced nature of physiological data and its influence on metrics and model outcomes. For example, in seizure detection settings, patient data consists mostly of interictal segments, with seizure events occurring rarely. Evaluations must reflect this natural imbalance to assess the model’s real-world utility faithfully.
@@
Standardized benchmarking tasks: The comparative analysis revealed significant heterogeneity in the tasks used for EEG-FM evaluation (see Table 2). The lack of common tasks across EEG-FM evaluations makes it challenging to understand the state of the art and highlights the need to identify a common core set of evaluations for future EEG-FM development. This set must cover multiple task types and include both classification and regression tasks, with dense (one label per EEG recording) and sparse (one label per EEG segment) labels. Additionally, the tasks must be challenging enough for previous generations of EEG-DL models, with ample room for improvement, unlike TUAB, where performance may already have saturated (85-87% accuracy) with traditional approaches [kiessner2024reaching]. Finally, the tasks must reflect the data heterogeneity encountered in clinical settings, including wake and sleep patient states.
@@
Trustworthy modeling: Despite the importance of explainability and uncertainty handling in the high-risk, expert-centric domain of medicine, experiments probing trust-related aspects of EEG-FMs have been limited. Notably, preliminary interpretability analyses based on intrinsic dimensionality and codebooks were conducted for BrainBERT and EEGFormer, respectively. However, studies that further demystify the EEG-FM black box are needed to gain insight into the knowledge learned by EEG-FMs (EEG patterns, dependencies, relationships) and the practical robustness of their decision-making process for downstream applications. Connections to known patterns of brain physiology or pathology may be necessary to make EEG-FMs trustworthy in the eyes of expert and clinical users.
@@ 6 Proposed Future Directions
From a time series perspective, EEG signals are distinct from general signals or sequences due to their non-linearity [klonowski2009everything], non-stationarity [klonowski2009everything], complex spatial relationships, and 1/f spectral characteristics [donoghue2020parameterizing]. These peculiarities would suggest that foundation modeling for EEG signals requires domain-specific inductive biases. Indeed, we have cultivated an EEG data-centric and domain-sensitive perspective throughout this review. However, counter-intuitively, EEG-FMs have enjoyed some success in drawing from vision and language domains in their adoption of patch-based sequences and transformer architecture design (transformer [vaswani2017attention], vision transformer [dosovitskiy2020image]), masked reconstruction-based SSL (masked autoencoders [he2022masked]), and discrete neural codebooks (VQ-VAE [van2017neural], BEiT v2 [peng2022beit]). Moreover, time series FMs pretrained on generic, non-neural time series have performed similarly to EEG-FMs on certain EEG tasks. Therefore, it is likely that future EEG-FMs could benefit from embracing both domain-specific insights and innovations from external data domains. These domains may be semantically distant to EEG, such as vision and natural language, or loosely related, such as non-neural time series, audio, speech, or other biosignals. Below, we outline the future research directions that we believe could support meaningful and sustained progress in the EEG-FM research domain. As illustrated in Figure 5, these recommendations are organized under three broad themes of development spanning benchmarks and software tools, technical advances, and real-world applications. These suggestions build upon the research gaps identified earlier and provide additional guidance for advancing and accelerating EEG-FM research and adoption.
Figure 5: Suggested future directions. (I) Benchmarks and tools: future EEG foundation models can be compared using standardized benchmarks against prevailing feature paradigms and participate in community-specific EEG challenges to establish their real-world utility. Frictionless and user-friendly software tools are needed to quickly adopt and experiment with off-the-shelf models. (II) Technical modeling: holistic evaluation frameworks that test embedding space semantics, robustness, and transfer efficiency can meaningfully track the state of the art. Advanced representation learning techniques, such as federated or multi-modal learning, can enhance large-scale pretraining. (III) Applications: collaborations with domain experts can inspire novel applications. Strategies that help identify suitable off-the-shelf models for a particular task and address translational hurdles, such as clinical interpretability, prospective validation, and operational feasibility, can increase adoption and impact.